OpenStudy (chrisneedshelp):

One of your classmates was absent on the day that you learned how to write and solve an equation from a contextual situation. Your teacher has asked you to explain the four steps and their application to the problem that follows. Natasha stole three fewer bases than twice the number Elena stole. If Natasha stole 15 bases, how many did Elena steal? SHOW YOUR WORK PLEASE :)

OpenStudy (anonymous):

good stuff

OpenStudy (chrisneedshelp):

Thanks :]

OpenStudy (anonymous):

et b be the number of bases. Natasha(N) Elena(E) 2E-3=15 2E=15+3 2E=18 E=18/2 E=9 My bad.

OpenStudy (chrisneedshelp):

not you mia dee

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

huuuwhhhee

OpenStudy (chrisneedshelp):

best response lol MIADEE YOUR A NUT GO SOMEHWERE